# Optimizing Token Choice for Code Watermarking: A RL Approach

### v1

@@ 4.2 How to Learn the Watermarked Policy?
As shown in Figure 1, to train the watermark model $\pi_{\phi}(a|\mathbf{c})$ in the watermarked policy $\pi_{\theta\oplus\phi}$, we employ GRPO as our RL framework. We selected GRPO for its computational efficiency and demonstrated effectiveness when integrating rule-based rewards on coding and mathematical tasks [7]. In a standard GRPO training scheme, three components are essential: a training policy, a reference policy, and reward functions. Our training pipeline uses $\pi_{\theta\oplus\phi}$ as the training policy rather than $\pi_{\phi}$ alone, since the complete generation process requires the coordinated operation of both the LLM and the watermark model. Importantly, we freeze the LLM parameters $\theta$ during training and optimize only watermark model parameters $\phi$, ensuring a portable and efficient watermark solution.
@@ 4.3 Policy Design
The implementation of our watermarked policy $\pi_{\theta\oplus\phi}$ presents fundamental technical challenges due to the non-differentiable nature of several components in Equation (1). Specifically, we implement the watermark model $\pi_{\phi}(a|\mathbf{c})$ as a transformer-based architecture. While the transformer outputs are continuous, we need to convert them to discrete outputs while maintaining differentiability. The watermark binary decision output $w$ and green token list assignment $G$ involve discrete selection processes that create optimization barriers in the neural network training pipeline.
@@
To enable gradient flow through the discrete decision processes for $w$, we apply straight-through estimation techniques. The watermark model transformer architecture first outputs a $|l_{\phi}|+1$ dimensional vector, which we represent as $(w_{\phi},l_{\phi})$. To derive our desired output $a=(w,G)$ from $(w_{\phi},l_{\phi})$, we implement the straight-through estimations:
$w=\mathbbm{1}_{w_{\phi}>0.5}+\sigma(w_{\phi})-\text{sg}(\sigma(w_{\phi})).$ (3)
During the forward pass, this formulation behave like the discrete operations $\mathbbm{1}[w_{\phi}>0]$. However, during backpropagation, gradients flow through the continuous relaxations $\sigma(w_{\phi})$, with $\text{sg}(\cdot)$ ensuring that only the gradient of the relaxation (and not their forward values) affects the computation. This technique enables end-to-end training despite the discrete nature of watermarking decisions.
@@
$\mathbf{g}=\mathbf{l}_{\phi}+(-\log(-\log(\mathbf{u})))$ (4)
@@
$\mathbf{l}_{G}=\mathbbm{1}_{v\in G}+\mathcal{S}(\mathbf{g})_{v}-\text{sg}(\mathcal{S}(\mathbf{g})_{v})$ (6)
where $\mathcal{S}(\mathbf{g})_{v}$ represents the Gumbel-Softmax-based continuous relaxation of the top-$k$ selection for token $v$, and $\text{sg}(\cdot)$ is the stop-gradient operator. During forward passes, we use hard discrete token selection based on token rankings from the Gumbel-perturbed logits, while during backward passes, the gradients flow through the softmax-based continuous relaxation. This dual-path approach allows the model to learn effective watermarking strategies while maintaining the discrete token selection.
For the reference policy component in our GRPO framework, we utilize the same architecture as the training policy $\pi_{\theta\oplus\phi}$, establishing a self-referential optimization approach. This design choice enables the watermarking parameters to be refined against their own previous distributions. The approach facilitates gradual refinement of watermark policy while ensuring that successive iterations remain coherent with previous behaviors, ultimately leading to more stable convergence and better performance.
@@ 4.4 Reward Function Design
The goal of the reward system is to leverage watermark detection feedback and code execution signals to guide watermarked policy $\pi_{\theta\oplus\phi}$ in selecting suitable watermarking positions ($w$) and making effective token choices for the green list ($G$). To achieve this, we have designed a reward system that comprises a set of reward functions balancing the objectives crucial for code watermarking.
Outcome-based Execution Reward $R_{1}$.
Our execution reward $R_{1}$ implements a binary evaluation of the watermarked code’s functional integrity through test case performance. We define:
[…]
@@
This straightforward formulation provides clear feedback on code correctness: successful execution of all test cases yields a positive reward, while any compilation error or runtime failure incurs a substantial penalty. By incorporating this binary signal into our reward function, we establish a strict constraint that compels the watermark policy to preserve the code’s original functionality.
Outcome-based Watermark Reward $R_{2}$.
@@
The watermark detection reward $R_{2}$ quantifies the statistical detectability of embedded watermarks within generated code sequences using a saturated function based on the statistical z-score. For each complete sequence $s$, we implement:
$R_{2}(s)=\begin{cases}1,&\text{if }z(s)\geq 3\\ \frac{z(s)}{3},&\text{if }0<z(s)<3\\ 0,&\text{if }z(s)\leq 0\end{cases}$ (8)
@@
where $z(s)$ represents the statistical z-score from our detection algorithm. This saturated reward formulation establishes clear boundaries: a z-score of 3 or higher (indicating strong statistical significance) receives the maximum reward of 1, while non-positive z-scores receive no reward. Between these thresholds, the reward scales linearly with the z-score. This design incentivizes the watermark policy to achieve statistically significant detection levels without wasting computational resources on unnecessarily strong signals.
Process-based Watermark Reward $R_{3}$.
@@
In our implementation, we find that outcome-based rewards alone are insufficient for effective policy learning, as they assign equal advantage to both red and green tokens when the overall sequence succeeds. To provide more granular guidance, we introduce a process-based reward $R_{3}$ that evaluates each token individually, offering immediate feedback on token selection decisions. This approach is straightforward to implement as we can directly check whether each generated token $s_{t}$ is green, red, or non-watermarked based on the policy’s action $a_{t}=(w_{t},G_{t})$ at the $t$-th token:
[…]
@@
This simple yet effective reward structure explicitly encourages the selection of green tokens during watermarking while penalizing red token selection. Non-watermarked positions receive a neutral reward, focusing the learning signal on positions where watermarking is applied. By providing direct feedback at each token generation step, the process-based reward creates a much clearer learning landscape compared to sequence-level outcome rewards alone. This token-by-token guidance significantly accelerates policy convergence and leads to more effective watermarking strategies.
Table 1: Performance (%) comparison of text detection methods. Methods marked with work in an impractical scenario with original LLM and prompts. Bold values indicate best performance within each category.
@@
Method / Metric Pass@1 Pass@10 AUROC TPR Pass@1 Pass@10 AUROC TPR
@@
Oracle Watermarks SWEET 61.77 75.45 83.10 42.68 41.01 48.35 84.40 32.00
Practical Watermarks WLLM 58.05 70.35 70.17 20.73 39.66 47.22 76.44 27.80
@@
CodeTracer 60.82 79.72 82.95 46.34 41.45 51.15 79.69 22.60
Integration into Advantage Computation.
Our approach calculates advantages at two complementary levels to optimize the watermarked policy. First, we compute outcome advantages $A_{1}$ from the execution and watermark detection rewards ($R_{1}$ and $R_{2}$) using GRPO’s outcome reward methodology. Second, we derive process advantages $A_{2}$ from the token-level watermark reward ($R_{3}$), normalizing these values across all tokens in all rollouts for each prompt. Our framework categorizes tokens into four distinct types: green tokens (watermarked and included in set $G$), red tokens (watermarked but falling in set $R$), non-watermarked tokens (where decision $w=0$), and non-code text tokens. To maintain focus specifically on code watermarking, we exclude non-code tokens entirely from process advantage calculations.
The final advantage function for each token $s_{t}$ integrates these signals directly:
$A_{\text{total}}(s_{t},a_{t})=A_{1}+A_{2}(s_{t},a_{t}),$ (10)
where $A_{1}$ represents outcome-based advantage and $A_{2}$ process-based advantage for token $s_{t}$ with action $a_{t}$. We then apply a binary mask to eliminate advantage values for any non-code tokens:
This integrated advantage function balances global sequence-level feedback with targeted token-specific guidance. While the outcome component $A_{1}$ evaluates tokens within the context of overall watermark effectiveness and functional code execution, the process component $A_{2}$ provides immediate direction for token selection decisions.
Ideally, our integrated reward system assigns high advantage to tokens that simultaneously maintain code functionality, contribute to effective watermark detection, and are selected from the green set. This comprehensive approach guides our watermarked policy $\pi_{\theta\oplus\phi}$ to identify optimal watermarking positions and select tokens that preserve functionality while introducing subtle, statistically detectable deviations from normal token distributions. By selectively masking non-code tokens, we concentrate the learning signal exclusively on elements most relevant to code watermarking, substantially improving policy optimization efficiency.
KL Regularization.
To ensure that the watermarked policy $\pi_{\theta\oplus\phi}$ does not deviate excessively from natural code generation patterns, thereby maintaining code stealth and linguistic quality, we incorporate a Kullback-Leibler (KL) divergence penalty into the overall GRPO loss function. This term, $\beta\cdot D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}})$, regularizes the watermarked policy against the reference policy $\pi_{\text{ref}}$.
5 Experiments
Evaluation Dimensions and Metrics.
We evaluate CodeTracer across three critical dimensions: (i) code functionality, (ii) watermark detectability, and (iii) robustness against attacks. For code functionality, we employ the Pass@k metric [3]. Watermark detection performance is measured by AUROC and TPR@5%FPR metrics.
Comparative Baselines.
For a comprehensive comparison, we compare CodeTracer against established post-hoc detection methods including logp(x), LogRank [5], DetectGPT [18], and GPTZero [21]. Moreover, we compare with active watermarking approaches such as WLLM [9] and EXP-edit [11]. We also include SWEET [13] serves as a reference baseline. However, to be noted, SWEET needs to redo the forward during the detection phase, which adopts a different setting from other baselines. A detailed discussion can be found in Appendix C.2.
Technical Implementation.
We primarily utilize OpenCoder-1.5B-Instruct [Huang2024OpenCoderTO] as the backbone LLM, with additional validation on the larger OpenCoder-8B-Instruct model. Performance is assessed on HumanEval [3] and MBPP [2] for Python code generation. Detailed implementation specifications, including dataset details, baseline configurations, model architectures, and training details are provided in Appendix C. In accordance with established reinforcement learning practices, we initially apply supervised fine-tuning to $\pi_{\phi}$ using proxy signals to establish fundamental code distribution knowledge through a combination of next token prediction and entropy distribution with binary cross entropy loss. Details and other techniques involved can be found in Appendix B.
5.1 Results on Watermark Detection Performance and Code Functionality
We present experiments evaluating CodeTracer’s performance on code functionality and watermark detection with HumanEval and MBPP in Table 1. CodeTracer outperforms other practical watermarking methods across most metrics. It shows minimal code quality degradation (Pass@1 scores of $60.82\%$ on HumanEval and $41.45\%$ on MBPP), while enabling effective detection (AUROC values of $82.95\%$ and $79.69\%$). With the highest TPR ($46.34\%$) on HumanEval among practical watermarks and detection performance comparable to oracle methods, CodeTracer emerges as the leading practical solution on these two datasets.
5.2 Robustness Analysis

### v2

@@ 4.1 How to Learn the Watermarked Policy?
As shown in Figure 1, we employ the GRPO algorithm (Shao et al., 2024) as our RL framework to train the watermark model $\pi_{\phi}(a|\mathbf{c})$ within the watermarked policy $\pi_{\theta\oplus\phi}$. GRPO offers computational efficiency and has demonstrated effectiveness in integrating rule-based rewards for coding and mathematical tasks (Guo et al., 2025). Our training pipeline requires three key components from the standard GRPO framework: a training policy, a reference policy, and reward functions. We use the integrated watermarked policy $\pi_{\theta\oplus\phi}$ as the training policy rather than solely $\pi_{\phi}$, as the complete code generation process requires coordinated operation of both the LLM and watermark model. Crucially, we freeze the LLM parameters $\theta$ during training and optimize only the watermark model parameters $\phi$. This design ensures a portable and efficient watermarking solution that trained watermark model can be applied to other pre-trained LLMs without requiring fine-tuning or additional training data.
4.2 Why Use RL for Learning?
Code watermarking presents a fundamental training paradox that reinforcement learning uniquely addresses. Traditional supervised learning requires pre-existing watermarked examples to train a detector, but creating such examples necessitates an already-trained watermark model: a circular dependency that renders supervised approaches infeasible. Reinforcement learning eliminates this paradox by learning through interaction rather than predefined examples.
Code watermarking is naturally suited for RL due to two key characteristics. First, watermark detection provides unambiguous, token-level verification signals that can be efficiently computed. Second, code functionality offers clear binary feedback through test case execution. Our RL formulation leverages these signals to offer distinct advantages: (1) it implicitly learns complex syntactic constraints across diverse programming languages without exhaustive manual specification; (2) it optimizes the watermarked policy to balance detectability and functionality through quantifiable reward mechanisms; (3) it enables end-to-end training without requiring pre-watermarked data.
@@ 4.3 Policy Design
The implementation of our watermarked policy $\pi_{\theta\oplus\phi}$ presents fundamental technical challenges due to the discrete nature of the watermark components $w$ and $G$ in Equation equation 1. We implement the watermark model $\pi_{\phi}(a|\mathbf{c})$ where $a=(w,G)$ as a transformer that outputs continuous values, but we need discrete outputs for watermarking decisions. Both the binary watermark decision $w$ and the green token list $G$ require discrete selections that impede gradient flow during training.
@@
To enable gradient flow through the discrete decision processes for watermark selection $w$, we apply straight-through estimation techniques. Specifically, our watermark transformer outputs a $(|\mathcal{V}|+1)$-dimensional vector as $(w_{\phi},l_{\phi})$, where $w_{\phi}\in\mathbb{R}$ is a scalar for watermark placement probability of and $l_{\phi}\in\mathbb{R}^{|\mathcal{V}|}$ as logit bias vector. To derive our desired output $a=(w,G)$ from $(w_{\phi},l_{\phi})$, we implement straight-through estimation:
$w=\mathbbm{1}_{w_{\phi}>0}+\sigma(w_{\phi})-\text{sg}(\sigma(w_{\phi})),$ (3)
where $\sigma(\cdot)$ is the sigmoid function and $\text{sg}(\cdot)$ is the stop-gradient operator. During the forward pass, this formulation behaves like the discrete operation $\mathbbm{1}_{w_{\phi}>0}$. However, during backpropagation, gradients flow through the continuous relaxation $\sigma(w_{\phi})$. The stop-gradient operator ensures only the gradient of the relaxation (and not its forward values) affects the computation. This technique enables end-to-end training despite the discrete nature of watermarking decisions made by $w$.
@@
$\mathbf{g}=\mathbf{l}_{\phi}+(-\log(-\log(\mathbf{u}))),$ (4)
@@
$\mathbf{l}_{G}=\mathbbm{1}_{v\in G}+\mathcal{S}(\mathbf{g})-\text{sg}(\mathcal{S}(\mathbf{g})),$ (6)
where $\mathcal{S}(\mathbf{g})$ represents the Gumbel-Softmax-based continuous relaxation of the top-$k$ selection, and $\text{sg}(\cdot)$ is the stop-gradient operator. During forward passes, we use hard discrete token selection $\mathbbm{1}_{v\in G}$ with Equation 5 based on token rankings from the Gumbel-perturbed logits $\mathbf{g}$, while during backward passes, the gradients flow through the softmax-based continuous relaxation $\mathcal{S}(\mathbf{g})$. This approach allows the model to learn effective watermarking while maintaining the discrete token selection.
For the reference policy in our GRPO framework, we utilize the same architecture as the training policy $\pi_{\theta\oplus\phi}$, establishing a self-referential optimization approach. This design choice enables the watermarking parameters to be refined against their own previous distributions. The approach facilitates gradual refinement of watermarked policy while ensuring that successive iterations remain coherent with previous behaviors, ultimately leading to more stable convergence and better performance.
@@ 4.4 Reward Function Design
Our reward system guides the watermarked policy $\pi_{\theta\oplus\phi}$ in watermark placement $w$ and green token selection $G$ using feedback from watermark detection and code execution. We design a multifaceted reward system that balances code functionality and watermark detectability.
Outcome-Based Execution Reward $R_{1}$. Our execution reward $R_{1}$ evaluates code functional correctness through binary assessment based on the performance on passing test cases:
$R_{1}=\begin{cases}1,&\text{if all test cases pass}\\ 0,&\text{otherwise}\end{cases}$ (7)
@@
This binary formulation provides unambiguous feedback on code correctness, where successful execution across all test cases yields a positive reward, while any compilation error, runtime failure, or test failure results in zero reward. By incorporating this strict binary signal into our reward function, we enforce functional preservation as a hard constraint, ensuring that the watermarked policy maintains code correctness while learning watermarking strategies.
@@
Outcome-Based Watermark Reward $R_{2}$. The watermark detection reward $R_{2}$ quantifies the statistical detectability of the watermark by employing a saturated function of the statistical $z$-score:
$R_{2}(s)=\begin{cases}1,&\text{if }z(s)\geq 4\\ \frac{z(s)}{4},&\text{if }0<z(s)<4\\ 0,&\text{if }z(s)\leq 0\end{cases}$ (8)
@@
where $z(s)$ is the statistical $z$-score obtained from the detection algorithm as a holistic measurement of watermark performance. This saturated reward formulation establishes clear boundaries: a $z$-score of 3 or higher (indicating strong statistical significance) receives the maximum reward of 1, while non-positive $z$-scores receive no reward. Between these thresholds, the reward scales linearly with the $z$-score. This design incentivizes our policy to achieve statistically significant detection levels.
@@
Process-Based Watermark Reward $R_{3}$. We observed that relying solely on outcome-based rewards can be insufficient for effective policy learning, as they provide the identical reward for tokens in a sequence. To provide more granular guidance, we introduce a process-based reward $R_{3}$ that evaluates individual generated tokens, offering immediate feedback on token selection decisions. This reward is straightforward to implement, as we can directly assess whether each generated token $s_{t}$ is green, red, or non-watermarked based on the policy’s action $a_{t}=(w_{t},G_{t})$ at time $t$:
$R_{3}(s_{t},a_{t})=\begin{cases}1,&\text{if }w_{t}=1\text{ and }s_{t}\in G_{t}\\ -1,&\text{if }w_{t}=1\text{ and }s_{t}\in R_{t}\\ 0,&\text{if }w_{t}=0\end{cases}$ (9)
@@
This reward structure explicitly encourages green token selection during watermarking while penalizing red token selection. Non-watermarked positions receive neutral rewards, concentrating the learning signal on watermarked tokens. By providing immediate feedback at each generation step, $R_{3}$ delivers more granular guidance than sequence-level rewards alone, significantly accelerating policy convergence and improving watermarking performance through direct token-level supervision.
Integration into Advantage Computation. Our approach calculates advantages at two complementary levels to optimize the watermarked policy. First, we compute outcome advantages $A_{1}$ from the execution reward $R_{1}$ and watermark detection reward $R_{2}$ using GRPO’s outcome-based group normalization mechanism. Second, we derive process advantages $A_{2}$ from the token-level watermark reward $R_{3}$, normalizing these values across all tokens in all rollouts for each prompt. We categorize generated tokens into four distinct types: green tokens (watermarked and belonging to set $G$), red tokens (watermarked but falling in set $R$), non-watermarked tokens (where $w=0$), and non-code text tokens. To maintain a specific focus on code watermarking, we exclude non-code tokens entirely.
The final advantage function for each token $s_{t}$ at time $t$ integrates these signals directly:
$A_{\text{total}}(s_{t},a_{t})=A_{1}+A_{2}(s_{t},a_{t}),$ (10)
where $A_{1}$ represents outcome-based advantage and $A_{2}$ represents process-based advantage for token $s_{t}$ with action $a_{t}$. We then apply a binary mask to eliminate advantage values for non-code tokens:
This integrated advantage function balances sequence-level feedback with token-specific guidance. $A_{1}$ provides sequence-level guidance from watermark effectiveness and functional correctness, while $A_{2}$ delivers immediate, token-level feedback for watermark positioning and green token selection.
Intuitively, our integrated reward system assigns high advantages to tokens that simultaneously maintain code functionality, contribute to effective watermark detection, and are selected from the green set. This approach guides our watermarked policy to identify optimal positions and select tokens that preserve functionality while introducing subtle, statistically detectable deviations.
KL Regularization. To ensure that the watermarked policy $\pi_{\theta\oplus\phi}$ does not deviate excessively from natural code generation patterns, thereby maintaining code stealth and linguistic quality, we incorporate a Kullback-Leibler (KL) divergence penalty into the overall GRPO loss function. This term, $D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}})$, regularizes the watermarked policy against the reference policy $\pi_{\text{ref}}$.
Objective of GRPO.
Combining all components discussed above, the final GRPO objective for our watermarked policy can be concisely expressed as:
$\max_{\phi}$ $\mathbb{E}_{s\sim\mathcal{D}}\Bigg[\frac{1}{|s|}\sum_{t=1}^{|s|}\min\Bigg[\frac{\pi_{\theta\oplus\phi}(s_{t}|s_{<t})}{\pi_{\text{ref}}(s_{t}|s_{<t})}\hat{A}(s_{t},a_{t}),\hskip-1.99997pt\text{clip}\left(\frac{\pi_{\theta\oplus\phi}(s_{t}|s_{<t})}{\pi_{\text{ref}}(s_{t}|s_{<t})},1-\varepsilon,1+\varepsilon\right)\hat{A}(s_{t},a_{t})\Bigg]\Bigg]$
$\hskip-1.99997pt-\beta D_{\text{KL}}(\pi_{\theta\oplus\phi}\parallel\pi_{\text{ref}}),$ (12)
where $\mathcal{D}$ represents rollout data from watermarked policy, $\hat{A}(s_{t},a_{t})$ is our masked advantage function incorporating both outcome-based and process-based rewards, $\varepsilon$ is a clipping parameter that limits policy updates, and $\beta$ controls KL regularization. This objective balances watermark detectability, code functionality, and natural generation patterns while only optimizing watermark parameters $\phi$.
5 Experiments
Evaluation Overview. We evaluate CodeTracer across five critical dimensions: (i) code functionality, (ii) watermark detectability, (iii) robustness against attacks, (iv) efficiency and (v) transferability. Following Lee et al. (2023), we employ two primary metrics: Pass@k for functionality assessment and AUROC and TPR@5%FPR for detection performance measurement (Chen et al., 2021).
Models and Datasets. We employ OpenCoder-1.5B-Instruct (Huang et al., 2024) as our primary backbone LLM and validate our findings on OpenCoder-8B-Instruct. We evaluate performance on three established benchmarks: HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) for Python evaluation, and HumanEvalPack (Muennighoff et al., 2023) for cross-language evaluation.
Baselines. We compare CodeTracer against two categories of baselines. For post-hoc detection methods, we evaluate against logp(x), LogRank (Gehrmann et al., 2019), DetectGPT (Mitchell et al., 2023), and GPTZero (Tian et al., 2023). For active watermarking approaches, we compare with WLLM (Kirchenbauer et al., 2023a), EXP-edit (Kuditipudi et al., 2023) and SWEET (Lee et al., 2023).
Technical Implementation. Detailed implementation specifications, including dataset details, baseline configurations, model architectures, and training details are provided in Appendix C. We mainly follow the experiment settings of Lee et al. (2023) to adapt it to our setting. In accordance with established reinforcement learning practices, we initially apply supervised fine-tuning to $\pi_{\phi}$ using proxy signals to establish code distribution knowledge through next token prediction. The training process only needs to be performed once and can be completed in 1 day on a single A100 GPU.
Table 1: Performance (%) comparison of detection methods. Bold: best performance for watermarks.
@@ 5 Experiments
Method Pass@1 Pass@10 AUROC TPR Pass@1 Pass@10 AUROC TPR
@@
Watermarks WLLM 58.05 70.35 70.17 20.73 39.66 47.22 76.44 27.80
@@
[…]
CodeTracer 62.65 77.11 77.71 32.32 42.10 52.01 78.42 31.60
Figure 2: Training on 1.5B LLM and evaluating on 8B OpenCoder-8B-Instruct LLM.
Table 2: Relative computational overhead of CodeTracer with a 1.5B base LLM. Our watermark model runs in parallel, so inference delay is only tensor slicing operation delay.
Component Parameters Memory Delay
[…]
1.5B LLM $\pi_{\theta}$ 1,500M  3-6GB  500-800ms
Relative Overhead < 10% < 15% < 0.02%
5.1 Main Results on Benchmark Datasets.
We present experiments evaluating CodeTracer’s performance on code functionality and watermark detection with HumanEval and MBPP in Table 1. We observe a clear improvement over other baselines.
Post-hoc vs. Watermarks. Post-hoc detection methods, including logp(x), LogRank, DetectGPT, and GPTZero, preserve original generation but demonstrate consistently poor discrimination performance. Their AUROC values range from 47.59% to 52.00% on HumanEval and 41.10% to 48.76% on MBPP, close to random guessing (50%). This indicates that with the advancement of LLMs, post-hoc detection methods struggle to distinguish AI-generated code from human-written code.
Code Functionality and Watermark Detectability. Watermarking methods demonstrate superior detection performance compared to post-hoc approaches, achieving AUROC values of 66.50-77.71% versus 47.59-52.00%. However, existing watermarking baselines suffer from noticeable code quality degradation, with Pass@1 scores dropping to 58.05-61.65% on HumanEval, representing a substantial trade-off between detection capability and functionality. CodeTracer achieves considerably better balance in this trade-off with the highest Pass@1 scores (62.65% on HumanEval, 42.10% on MBPP) and superior detection capability (AUROC of 77.71% and 78.42%), consistently outperforming SWEET and other baselines across both functionality and watermark detectability metrics.
[…]
00footnotetext: SWEET and SWEET use different entropy thresholds to explore detectability-functionality trade-offs. We adapt SWEET to use fixed context windows (vs. full generation history and original LLM) for practical and fair comparison with other baselines. See Appendix C.2 for details about entropy threshold setup.